OpenStudy (twdlover23):

A car traveling at 26 m/s starts to decelerate steadily it comes to a complete stop in 7 seconds. What is its acceleration?

OpenStudy (icrazyaliens):

First, Vi= 26 m/s ( initial Velocity) Vf= 0 m/s ( Final Velocity) because it comes to stop t= 7 s a= ? the acceleration must be in a negative sign because it is decelerated second, \[Xf-X0=Vi \times t + (1/2) \times a \times t^2\] solve the equation for a \[a=2( \Delta x - Vi \times t )/t^2\] add this equation to this \[Vf = Vi + a \times t \] then solve for x and find the position once you find the position solve for the acceleration by this formula \[Vf^2 = Vi^2 + 2a \Delta x\]minus sign indicates that the acceleration is decreased. The answer is \[a = -3.71 m/s^2\]

OpenStudy (twdlover23):

thank you so much! I understand now